# Marcus Webb

### v2

@@ 1 Introduction
where $1\leq p\leq\infty$, and $Z_{n}(L_{\mu}^{2}(E),L_{\nu}^{p}(F))$ is the Cauchy–Zolotarev number,
@@ 2 Proof of the Main Theorem
$\mathcal{K}[g](x)=\frac{1}{2\pi\mathrm{i}}\int_{E}\int_{\Gamma}\frac{K(x,\xi)}{y-\xi}\,\mathrm{d}\xi\,g(y)\,\mathrm{d}\mu(y)=\frac{1}{2\pi\mathrm{i}}\int_{\Gamma}K(x,\xi)\,\left(\int_{E}\frac{g(y)}{y-\xi}\,\mathrm{d}\mu(y)\right)\,\mathrm{d}\xi.$ (2.2)
@@ 3 Cauchy–Zolotarev numbers
$Z_{n}([a,b],[c,d])\leq 4\exp\left(\frac{-n\pi^{2}}{\log(16\gamma)}\right),\qquad\gamma=\frac{|c-a||d-b|}{|c-b||d-a|}.$
@@
Now we can substitute the form of $t$ into the logarithm. Since $y\to\log\left(\frac{y-b}{t-b}\right)$ is a monotonic function, its extrema are attained at the endpoints. For this choice of $t$, the endpoint values are $\pm\frac{1}{2}\log\left(\frac{d-b}{c-b}\right)$, which yields the desired bound on $Z_{1}$. To bound $Z_{n}$ for $n\geq 1$, use Lemma 3.2 with $s=1$. ∎
@@ 6.1 Cauchy matrices and tensors
Figure 3: Cauchy matrices and tensors. On the left we used $A\in\mathbb{R}^{N\times N}$ in Equation (6.1) with $N=100$, $D=[1,70]$, $E=[2,100]$. On the right we used $A\in\mathbb{R}^{N\times N\times N}$ in Equation (6.5) with $N=50$, $D=[1,70]\times[1,199]\subset\mathbb{C}^{2}$, $E=[2,100]$. In both plots the sample points are equispaced over their respective intervals. For the purple triangles in the right plot, I numerically calculated the optimal rational function for $Z_{n}(\{(w_{k},x_{i})\},\{y_{j}\})$ and evaluated the associated rational interpolant of $K$.
@@ 6.2 Log-Cauchy matrices
Figure 4: Log-Cauchy matrices. We used $A\in\mathbb{R}^{N\times N}$ in Equation (6.9) with $N=100$, $D=E=[1,N]$ and $\{x_{i}\}$, $\{y_{j}\}$ randomly uniformly distributed in $[1,N]$ (and $x_{1}=y_{1}=1$, $x_{N}=y_{N}=N$). The Chebyshev interpolant (blue dots) uses the modified interpolation points mentioned in Remark 5.4. The Zolotarev interpolant (yellow circles) uses the $\phi$ in the proof of Lemma 3.5 for the nodes and poles, whereas the suboptimal Zolotarev interpolants (cyan plus signs) uses the $\phi$ that is optimal for $Z_{n}(E,F)$.
@@ 6.3 Twisted Hankel transform matrix
$K(x,y)=\mathrm{H}_{0}^{(1)}(xy)\mathrm{e}^{-\mathrm{i}xy},\qquad x\in[0,1],\qquad y\in[\omega_{1},\omega_{N}],$ (6.17)
where $\mathrm{H}_{0}^{(1)}$ is the Hankel function of the first kind. For each $x\in[0,1]$, $K(x,\cdot)$ is analytic in $\mathbb{C}\setminus(-\infty,0]$, where $\mathrm{H}_{0}^{(1)}$ has its principle branch cut. Therefore, we can take $D=[0,1]$, $E=[\omega_{1},\omega_{N}]$, $F=[-\infty,0]$. Its Grothendieck dual $\mathcal{K}^{\prime}:H([-\infty,0])\to C([0,1])$ is
@@
by (NIST:DLMF, , 10.7.8). This second fact ensures that the contour integrals depending on $R$ will converge to zero as $R\to\infty$. Therefore, as a member of $H^{\prime}([-\infty,0];L_{\lambda}^{2}([0,1]))$, the Grothendieck dual has kernel $K^{\prime}(x,z)=\frac{2}{\pi\mathrm{i}}\mathrm{J}_{0}(xz)\mathrm{e}^{-\mathrm{i}xz}$, so we can use Lemma 4.2 to bound its norm
$\|\mathcal{K}^{\prime}\|_{H^{1}([-\infty,0])\to L_{\lambda}^{2}([0,1])}=\sup_{z\leq 0}\frac{2}{\pi}\left(\int_{0}^{1}|\mathrm{J}_{0}(xz)|^{2}\,\mathrm{d}\lambda(x)\right)^{1/2}=\frac{2}{\pi}\lambda([0,1])^{1/2},$ (6.20)
@@
$Z_{n}(L_{\mu}^{2}(\omega_{1},\omega_{N}),L^{1}(-\infty,0))\leq\frac{1}{2}\,\log\left(\frac{\omega_{N}}{\omega_{1}}\right)\,\mu([\omega_{1},\omega_{N}])^{1/2}\,Z_{n-1}([\omega_{1},\omega_{N}],[-\infty,0]).$ (6.21)
@@
$\|\mathcal{K}-\mathcal{K}_{n}\|_{L_{\mu}^{2}([\omega_{1},\omega_{N}])\to L_{\lambda}^{2}([0,1])}\leq\frac{N}{\pi}\,\log\left(\frac{\omega_{N}}{\omega_{1}}\right)\,Z_{n-1}([\omega_{1},\omega_{N}],[-\infty,0]).$ (6.22)
@@
$A_{i,j}=K(\omega_{i}/\omega_{N},\,\omega_{j})\qquad i,j=1,2,\ldots,N.$ (6.23)
The purple triangles correspond to the interpolation points and poles taken to be the roots and poles of the Zolotarev rational function for $Z_{n}(\{\omega_{1},\omega_{2},\ldots,\omega_{N}\},[-\infty,0])$. To compute this $\phi$, I first mapped these sets to bounded sets with the Möbius transformation $T(z)=\frac{1}{\omega_{N}+1}$. I used Trefethen and Wilber’s algorithm trefethen2025computation , which requires discrete sets as input, so I approximated $T([-\infty,0])$ by 2000 equally spaced points between 0 and $(1+\omega_{N})^{-1}$. Then I mapped the resulting zeros and poles (which are covariant under Möbius transformations) back to the original sets. The resulting interpolants (purple triangles) achieved a smaller error than those implied by Equation (6.22) (yellow circles).
I did not use the extra interpolation point $t$ from the proof of Lemma 3.5, because the results were, for an unknown reason, slightly worse. The results of this paper belie some difficulties in obtaining consistent results from Trefethen and Wilber’s algorithm (which the authors describe as “90% reliable” trefethen2025computation ).
@@
Figure 5: Twisted Hankel transform matrix. We used $A\in\mathbb{R}^{N\times N}$ in Equation (6.23) with $N=100$. The Chebyshev interpolant (blue dots) uses the modified interpolation points mentioned in Remark 5.4. The Zolotarev interpolant (yellow circles) uses the $\phi$ in the proof of Lemma 3.5 for the nodes and poles, whereas the semidiscrete Zolotarev interpolants (purple triangles) use the $\phi$ that is optimal for $Z_{n}(\{\omega_{1},\omega_{2},\ldots,\omega_{N}\},[-\infty,0])$ as described in the main text.
@@ 7 Conclusion
The choice of interpolation points and poles is ripe for investigation. More research on computation of Zolotarev rational functions associated with continuous and, more importantly, discrete subsets of $\widehat{\mathbb{C}}$ is needed, since Trefethen and Wilber’s algorithm is not 100% reliable trefethen2025computation . Figure 4 highlights the potential of direct computation of the optimal rational functions for Cauchy-Zolotarev numbers. When $F$ is unknown, the best approach to finding good points and poles may involve the AAA algorithm nakatsukasa2018aaa applied to $K$. Developing a practical interpolation scheme in this setting could have significant consequences, impacting any algorithm employing low-rank approximation of analytic kernels.
@@ Acknowledgements
The author thanks Andrew Horning, Daan Huybrechs, Jonas Latz, Yuji Nakatuskasa, Will Parnell, Endre Süli, Françoise Tisseur, Alex Townsend and Heather Wilber for helpful comments and discussion. Probing questions from Georg Maierhofer and Nick Trefethen led to improvements in the mathematics and the presentation.
@@
which tends to zero as $\varepsilon\to 0$ and then $R\to\infty$. ∎

### v3

@@ 1 Introduction
where $1\leq p\leq\infty$, the number $Z_{n}(L_{\mu}^{2}(E),L_{\nu}^{p}(F))$ is the Cauchy–Zolotarev number,
@@ 2 Proof of the Main Theorem
$\mathcal{K}[g](x)=\frac{1}{2\pi\mathrm{i}}\int_{E}\left(\int_{\Gamma}\frac{K(x,\xi)}{y-\xi}\,\mathrm{d}\xi\right)\,g(y)\,\mathrm{d}\mu(y)=\frac{1}{2\pi\mathrm{i}}\int_{\Gamma}K(x,\xi)\,\left(\int_{E}\frac{g(y)}{y-\xi}\,\mathrm{d}\mu(y)\right)\,\mathrm{d}\xi.$ (2.2)
@@ 3 Cauchy–Zolotarev numbers
$Z_{n}([a,b],[c,d])\leq 4\exp\left(\frac{-n\pi^{2}}{\log(16\gamma)}\right),\quad\text{ where }\quad\gamma=\frac{(c-a)(d-b)}{(c-b)(d-a)}.$
@@
Note that (db)/(cb)(d-b)/(c-b) is $\gamma$ in the limiting case $a\to$-$\infty$.
@@
Now we can substitute the form of $t$ into the logarithm. Since $y\mapsto\log\left(\frac{y-b}{t-b}\right)$ is a monotonic function, its extrema are attained at the endpoints. For this choice of $t$, the endpoint values are $\pm\frac{1}{2}\log\left(\frac{d-b}{c-b}\right)$, which yields the desired bound on $Z_{1}$. To bound $Z_{n}$ for $n>1$, use Lemma 3.2 with $s=1$. ∎
@@ 6.1 Cauchy matrices and tensors
Figure 3: Cauchy matrices and tensors. On the left I used $A\in\mathbb{R}^{N\times N}$ in Equation (6.1) with $N=100$, $D=[1,70]$, $E=[2,100]$. On the right I used $A\in\mathbb{R}^{N\times N\times N}$ in Equation (6.5) with $N=50$, $D=[1,70]\times[1,199]\subset\mathbb{C}^{2}$, $E=[2,100]$. In both plots the sample points are equispaced over their respective intervals. For the purple triangles in the right plot, I numerically calculated the optimal rational function for $Z_{n}(\{(w_{k},x_{i})\},\{y_{j}\})$ and evaluated the associated rational interpolant of $K$.
@@ 6.2 Log-Cauchy matrices
Figure 4: Log-Cauchy matrices. I used $A\in\mathbb{R}^{N\times N}$ in Equation (6.9) with $N=100$, $D=E=[1,N]$ and $\{x_{i}\}$, $\{y_{j}\}$ randomly uniformly distributed in $[1,N]$ (and $x_{1}=y_{1}=1$, $x_{N}=y_{N}=N$). The Chebyshev interpolant (blue dots) uses the modified interpolation points mentioned in Remark 5.4. The Zolotarev interpolant (yellow circles) uses the $\phi$ in the proof of Lemma 3.5 for the nodes and poles, whereas the suboptimal Zolotarev interpolants (cyan plus signs) uses the $\phi$ that is optimal for $Z_{n}(E,F)$.
@@ 6.3 Twisted Hankel transform matrix
$K(x,y)=\mathrm{H}_{0}^{(1)}(xy)\mathrm{e}^{-\mathrm{i}xy},\qquad x,y\in\left[\widehat{\omega}_{1},\widehat{\omega}_{N}\right],\qquad\widehat{\omega}_{j}=\omega_{j}/\sqrt{\omega_{N+1}},$ (6.17)
where $\mathrm{H}_{0}^{(1)}$ is the Hankel function of the first kind. For each $x\in[\widehat{\omega}_{1},\widehat{\omega}_{N}]$, $K(x,\cdot)$ is analytic in the set $\mathbb{C}\setminus(-\infty,0]$, where $\mathrm{H}_{0}^{(1)}$ has its principle branch cut. Therefore, we can take $D=E=[\widehat{\omega}_{1},\widehat{\omega}_{N}]$ and $F=[-\infty,0]$. Its Grothendieck dual $\mathcal{K}^{\prime}:H([-\infty,0])\to C([\widehat{\omega}_{1},\widehat{\omega}_{N}])$ is
@@
by (NIST:DLMF, , 10.7.8). This second fact ensures that the contour integrals depending on $R$ converge to zero as $R\to\infty$. Therefore, as a member of $H^{\prime}([-\infty,0];L_{\lambda}^{2}([\widehat{\omega}_{1},\widehat{\omega}_{N}]))$, the Grothendieck dual has kernel $K^{\prime}(x,z)=\frac{2}{\pi\mathrm{i}}\mathrm{J}_{0}(xz)\mathrm{e}^{-\mathrm{i}xz}$, so we can use Lemma 4.2 to bound its norm
$\|\mathcal{K}^{\prime}\|_{H^{1}([-\infty,0])\to L_{\lambda}^{2}([\widehat{\omega}_{1},\widehat{\omega}_{N}])}=\sup_{z\leq 0}\frac{2}{\pi}\left(\int_{\widehat{\omega}_{1}}^{\widehat{\omega}_{N}}|\mathrm{J}_{0}(xz)|^{2}\,\mathrm{d}\lambda(x)\right)^{1/2}=\frac{2}{\pi}\lambda([\widehat{\omega}_{1},\widehat{\omega}_{N}])^{1/2},$ (6.20)
@@
$Z_{n}(L_{\mu}^{2}([\widehat{\omega}_{1},\widehat{\omega}_{N}]),L^{1}(-\infty,0))\leq\frac{1}{2}\,\log\left(\frac{\omega_{N}}{\omega_{1}}\right)\,\mu([\widehat{\omega}_{1},\widehat{\omega}_{N}])^{1/2}\,Z_{n-1}\left([-\infty,0],\left[1,\frac{\omega_{N}}{\omega_{1}}\right]\right).$ (6.21)
@@
$\|\mathcal{K}-\mathcal{K}_{n}\|_{L_{\mu}^{2}([\widehat{\omega}_{1},\widehat{\omega}_{N}])\to L_{\lambda}^{2}([\widehat{\omega}_{1},\widehat{\omega}_{N}])}\leq\frac{N}{\pi}\,\log\left(\frac{\omega_{N}}{\omega_{1}}\right)\,Z_{n-1}\left([-\infty,0],\left[1,\frac{\omega_{N}}{\omega_{1}}\right]\right).$ (6.22)
@@
$A_{i,j}=K(\widehat{\omega}_{i},\,\widehat{\omega}_{j})\qquad i,j=1,2,\ldots,N.$ (6.23)
The purple triangles correspond to the interpolation points and poles taken to be the roots and poles of the Zolotarev rational function for $Z_{n}(\{\widehat{\omega}_{1},\widehat{\omega}_{2},\ldots,\widehat{\omega}_{N}\},[-\infty,0])$. To compute this $\phi$, I first mapped these sets to bounded sets with the Möbius transformation $T(z)=\frac{1}{1+\widehat{\omega}_{N}-z}$. I used Trefethen and Wilber’s algorithm trefethen2025computation , which requires discrete sets as input, so I approximated $T([-\infty,0])$ by 2000 equally spaced points between 0 and $(1+\widehat{\omega}_{N})^{-1}$. Then I mapped the resulting zeros and poles (which are covariant under Möbius transformations) back to the original sets. The resulting interpolants (purple triangles) achieved a smaller error than those implied by Equation (6.22) (yellow circles).
For the semidiscrete interpolant, I did not use the extra interpolation point $t$ from the proof of Lemma 3.5, because the results were, for an unknown reason, slightly worse. The results of this paper belie some difficulties in obtaining consistent results from Trefethen and Wilber’s algorithm (which the authors describe as “90% reliable” trefethen2025computation ).
@@
Figure 5: Twisted Hankel transform matrix. I used $A\in\mathbb{R}^{N\times N}$ in Equation (6.23) with $N=100$. The Chebyshev interpolant (blue dots) uses the modified interpolation points mentioned in Remark 5.4. The Zolotarev interpolant (yellow circles) uses the $\phi$ in the proof of Lemma 3.5 for the nodes and poles, whereas the semidiscrete Zolotarev interpolants (purple triangles) use the $\phi$ that is optimal for $Z_{n}(\{\widehat{\omega}_{1},\widehat{\omega}_{2},\ldots,\widehat{\omega}_{N}\},[-\infty,0])$ as described in the main text.
@@ 7 Conclusion
The choice of interpolation points and poles is ripe for investigation. More research on computation of Zolotarev rational functions associated with continuous and, more importantly, discrete subsets of $\widehat{\mathbb{C}}$ is needed, since Trefethen and Wilber’s algorithm is not 100% reliable trefethen2025computation . Figure 4 highlights the potential of direct computation of the optimal rational functions for Cauchy-Zolotarev numbers. When $F$ is unknown, the best approach to finding good points and poles may involve the AAA algorithm nakatsukasa2018aaa applied to $K$. Developing a practical interpolation scheme in this setting could have significant consequences, impacting any application employing low-rank approximation of analytic kernels.
@@ Acknowledgements
The author thanks Andrew Horning, Yanghong Huang, Daan Huybrechs, Jonas Latz, Yuji Nakatuskasa, Will Parnell, Endre Süli, Françoise Tisseur, Alex Townsend and Heather Wilber for helpful comments and discussion. Probing questions from Georg Maierhofer and Nick Trefethen led to improvements in the mathematics and the presentation.
@@
which tends to zero as $R\to\infty$. ∎